# Learning Individual Behavior in Agent-Based Models with Graph Diffusion Networks

### v1

@@ Abstract
Agent-Based Models (ABMs) are powerful tools for studying emergent properties in complex systems. In ABMs, agent behaviors are governed by local interactions and stochastic rules. However, these rules are, in general, non-differentiable, limiting the use of gradient-based methods for optimization, and thus integration with real-world data. We propose a novel framework to learn a differentiable surrogate of any ABM by observing its generated data. Our method combines diffusion models to capture behavioral stochasticity and graph neural networks to model agent interactions. Distinct from prior surrogate approaches, our method introduces a fundamental shift: rather than approximating system-level outputs, it models individual agent behavior directly, preserving the decentralized, bottom-up dynamics that define ABMs. We validate our approach on two ABMs (Schelling’s segregation model and a Predator-Prey ecosystem) showing that it replicates individual-level patterns and accurately forecasts emergent dynamics beyond training. Our results demonstrate the potential of combining diffusion models and graph learning for data-driven ABM simulation.
@@ 1 Introduction
Agent-Based Models (ABMs) are computational frameworks in which autonomous “agents” interact with each other and their environment, leading to emergent collective behavior (wilensky2015introduction, ). ABMs are typically characterized by: (i) a well-defined network of interactions, where the state of each agent is influenced by the states of a specific set of other agents, usually from the previous time step; (ii) stochasticity, meaning that agents’ decisions incorporate a degree of randomness, producing probability distributions over multiple runs that capture real-world uncertainty and variation. ABMs have proven to be a powerful tool for developing and refining theoretical understanding, particularly in identifying minimal sets of micro-level rules that generate realistic macro-level outcomes (railsback2019agent, ). In this sense, they have been instrumental in modeling a diverse range of phenomena (castellano2009statistical, ), including structure formation in biological systems, pedestrian traffic, urban aggregation, and opinion dynamics. More recently, ABMs have demonstrated their value as forecasting tools (poledna2020economic, ), such as in predicting the economic impacts of the COVID-19 pandemic (pichler2022forecasting, ).
@@
However, this progress is occurring despite the absence of principled methods to systematically align ABMs with real-world data. While various approaches have been proposed for calibrating macro-level parameters of ABMs (platt2020comparison, ), there are still no established methods for tuning the micro-level behaviors of individual agents to data. One potential approach is to manually construct a probabilistic model that replicates the ABM and then use its likelihood function to estimate individual state variables (monti2023learning, ). However, this method requires the manual development of an ad-hoc probabilistic framework that reproduces the original ABM. Thus, what is currently missing is a fully automated method for deriving a learnable, differentiable model directly from an ABM.
@@
Figure 1: Overview of the training and data generation pipeline for differentiable surrogates of Agent-Based Models. The top-left panel illustrates the training process: we run simulations using the original ABM, and use the resulting trajectories to train the differentiable surrogate. The top-right panel visualizes the structure of the ABM-generated data using the Predator-Prey model as an example: a state at time step $t{-}1$ gives rise to multiple possible states at time $t$, one of which is chosen to generate further possible states at $t{+}1$. Colored cells highlight the behavior of a specific “prey” agent — green for “move,” red for “die,” and pink for “reproduce.” The bottom panel shows the data generation phase: given a new observed state, the trained surrogate can simulate plausible future states, effectively mimicking the original ABM’s generative behavior.
@@
In this work, we propose a novel approach to address this challenge: combining a graph neural network with a diffusion model to learn a differentiable surrogate of an ABM, from ABM-generated data. We refer to this method as a Graph Diffusion Network (GDN). In this framework, a graph neural network captures the interactions that govern the evolution of each agent’s state in response to other agents, while the diffusion model learns the distribution of possible state transitions, conditioned on these interactions. A central aspect of our approach is its explicit modeling of individual agent behavior. Rather than treating the system as a whole, we focus on how each agent acts as an independent entity, while also incorporating the influence of other agents on its decisions. This approach ensures that the emergent dynamics remain faithful to the decentralized nature of ABMs.
@@
Our approach draws inspiration from previous work on using neural network models to emulate deterministic cellular automata (grattarola2021learning, ). However, we extend this idea to the broader domain of ABMs by introducing a crucial component: stochasticity. By incorporating stochasticity, our architecture can learn directly from ABM-generated data traces, making it adaptable to a wide variety of agent-based models across diverse real-world applications. Furthermore, since our method is trained on data traces, it can seamlessly integrate empirical observations alongside simulated data, thus being potentially applicable to real-world scenarios. In this sense, our work represents a first step toward developing a comprehensive methodology for creating easy-to-use, learnable ABMs.
@@ 2 Background
To illustrate this structure, we consider two ABMs used throughout the paper. The first is the well-known model by Schelling schelling1971dynamic , where $\mathbf{Z}_{t}$ captures agents positions and colors, and $\Theta$ indicates preference for same-color neighbors. Even with some tolerance for neighbors of different color, agents often form segregated clusters (ubarevivciene2024fifty, ). This clear mismatch between individual preferences and aggregate outcomes is a classic example of emergence. The second model is a predator-prey model (tang2022data, ; wilensky2015introduction, ), describing the ecological dynamics between two interacting species, with one acting as predator and the other as prey, similarly to the Lotka-Volterra equations. In this ABM, $\mathbf{Z}_{t}$ includes agent position and type (prey-predator), while $\Theta$ governs the probability to move, reproduce or die. This model replicates the cyclical predator-prey population dynamics, typical of Lotka-Volterra systems, while also capturing complex spatial patterns reminiscent of spatial evolutionary games (nowak1992evolutionary, ).
@@
ABMs have traditionally been powerful for theory generation, but in recent years, they have become increasingly data-driven (pangallo2024datadriven, ). To align ABM output with empirical data, most efforts focus on calibrating parameters $\Theta$ so that model-generated summary statistics match observed ones (platt2020comparison, ; quera2023blackbirds, ). Less attention, however, has been paid to estimating agent states $\mathbf{Z}_{t}$, which is key for matching time series further to summary statistics. Some researchers use data assimilation methods like particle filters (lux2018estimation, ) or ensemble Kalman filters (oswald2025agent, ) to infer $\mathbf{Z}_{t}$. A more principled alternative is to make ABMs differentiable, enabling the maximization of a likelihood function via gradient descent and automatic differentiation (monti2020learning, ; monti2023learning, ). While differentiability is straightforward for simple stochastic behaviors, such as those governed by Bernoulli trials (arya2022automatic, ), it becomes far more challenging for complex behaviors like those observed in the Schelling and predator-prey models.
@@
To address this and other challenges in ABMs, researchers have increasingly turned to more tractable surrogates, also known as meta-models or emulators (kennedy2001bayesian, ; farah2014bayesian, ; glielmo2023reinforcement, ). Surrogate models typically learn directly the mapping from parameters $\Theta$ to static summary statistics, disregarding individual behavior and model dynamics. For instance, a surrogate in Lamperti et al. lamperti2018agent maps $\Theta$ to the mean growth rate of the economy. More recent research has also explored the emulation of model dynamics. Grattarola et al. grattarola2021learning use Graph Neural Networks to approximate cellular automata, which can be seen as a special case of ABMs with deterministic interaction rules. Dyer et al. dyer2024interventionally propose Ordinary Differential Equation emulators to construct interventionally consistent surrogates, ensuring that micro-state interventions produce results aligned with macro-state interventions. Casert et al. casert2024learning employ Transformers to model the transition of physical systems from one configuration to another, effectively capturing the evolution of the distribution of micro-states over time, but modeling in terms of configuration transition rates rather than reproducing individual agent behavior.
@@
To achieve this goal, our framework relies on a novel combination of graph neural networks and diffusion models. Diffusion models (ddpm_original_paper, ) were first introduced in the context of image generation, where they demonstrated impressive generation capabilities (diffusion_vision_review, ), and were then applied to other domains (tabddpm, ). A number of works addressed graph data (graph_diffusion_review, ), for example in molecule modeling (molecule_modeling, ) and protein structure generation (protein_diffusion, ). However, these works focus on the generation of graphs, while our architecture learns to generate random samples that are conditioned on information found on a graph. To the best of our knowledge, our work is the first application of this generative framework to individual behavior modeling in simulation systems, such as ABMs.
@@ 3 Methods
To effectively model these components in the same individual-level view, we leverage respectively (i) message-passing GNNs (message_passing_gnn, ), which model the relationship between the evolution of an agent’s state and the state of its neighbors on the graph; (ii) conditional diffusion models (conditionalimagesynthesisdiffusionreview, ), generative architectures well-suited to learning complex, multimodal distributions, allowing us to capture the intrinsic stochasticity of agent behaviour.
@@
Overview. In order to learn the distribution $P_{\Theta}$, the training phase requires observations of different outcomes given the same starting conditions. To do so, in our framework we use the original ABM to generate a data set as a ramification of possible states, namely $\bigl{(}\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}\bigr{)}\;\longrightarrow\;\mathbf{Z}_{t+1}^{(i)}$ (see Figure 1). Our Graph Diffusion Network then approximates the stochastic kernel $P_{\Theta}$ by integrating a Message-Passing GNN with a Conditional Diffusion Model, of learnable parameters $\omega$ and $\phi$ respectively. The GNN aggregates each agent’s state $\mathbf{Z}_{t}^{(i)}$ and its neighbours’ states $\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}$ via permutation-invariant message and readout functions to produce an interaction embedding $\mathbf{g}_{t}^{(i)}$. This embedding acts as a compact representation of the information coming from $i$’s neighbors at $t$, affecting the distribution of possible states of agent $i$ at time $t+1$. As such, it is passed to the diffusion model: conditioned on $\mathbf{Z}_{t}^{(i)}$ and $\mathbf{g}_{t}^{(i)}$, the diffusion model learns to transform a sample of Gaussian noise into a possible instance of the next state $\mathbf{Z}_{t+1}^{(i)}~{}$. By minimizing the standard denoising loss over all observed transitions, this hybrid architecture captures both the graph-structured interactions and the inherent stochasticity of agent dynamics.The trained model $GDN_{\phi,\omega}$ is therefore able to generate, given a graph $G_{t}$ of interacting agents and the state of each one $\mathbf{Z}^{(i)}_{t}$, a sequence of possible next states $\mathbf{Z}^{(i)}_{t+1}$. The consecutive application of $GDN_{\phi,\omega}$ allows to reproduce the behavior of the original model. We now describe in detail each of these components.
@@
Message-passing GNN. The GNN operates on the provided interaction graph $G_{t}=(A,E_{t})$, that we assume to be known or to be computable from $\mathbf{Z}_{t}$ (e.g., in the Schelling model, the position of the agents determine who interacts with whom). For each agent $i$, the GNN aggregates its state $\mathbf{Z}_{t}^{(i)}$ together with each neighbor’s state $\mathbf{Z}_{t}^{(j)}$ via a permutation-invariant operator $\bigoplus$, and then feeds the concatenated result through an MLP $f_{\omega}$ (message_passing_gnn, ), that is $\mathbf{g}_{t}^{(i)}=f_{\omega}\!\Bigl{(}\mathbf{Z}_{t}^{(i)},\,\bigoplus_{j\in N_{t}^{(i)}}\bigl{(}\mathbf{Z}_{t}^{(i)},\mathbf{Z}_{t}^{(j)}\bigr{)}\Bigr{)}$. The resulting vector $\mathbf{g}_{t}^{(i)}$ captures how $i$’s local neighborhood influences its next-state distribution.
@@
Conditional diffusion model. The diffusion model then learns the distribution over future states given this output from the graph and the state of a given agent. Diffusion models do so by reversing a fixed Gaussian noising process (ddpm_original_paper, ). The obtained denoising process, indexed by $\tau\in\{\tau_{\text{max}},\dots,0\}$, starts from a sample of Gaussian noise $\mathbf{x}_{\tau_{max}}\sim\mathcal{N}(\mathbf{0,I})$ and in a sequence of denoising diffusion steps transforms it into a possible next state $\mathbf{x}_{0}\approx\mathbf{Z}_{t+1}^{(i)}$. In this setting, we denote the general latent $\mathbf{x}_{\tau}$ as $\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau)$. Each step of this process receives as input (i) the agent’s current state $\mathbf{Z}_{t}^{(i)}$, (ii) its interaction embedding $\mathbf{g}_{t}^{(i)}$, and (iii) a sinusoidal positional embedding of $\tau$ (attention_is_all_you_need, ). These inputs are first transformed by MLPs to form the condition vector $\mathbf{c}_{t}^{(i)}$. Then, a feed-forward network $\phi$ is trained to predict the noise residual $\epsilon_{\phi}$, i.e., the change to apply to the input $\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau)$ to continue the denoising process.
@@
Ramification data set. Given these two components, our framework uses the original ABM to produce a ramifications data set (see Figure 1). Such data set follows one main branch that specifies the evolution of the ABM, and multiple alternative stochastic evolutions of each time step from time $t$ to time $t+1$. This method makes it possible to expose the model to multiple stochastic successors from identical conditioning context, while avoiding exponential growth in the number of histories. Starting from an initial configuration $\mathbf{Z}_{0}=\bigl{(}\mathbf{Z}_{0}^{(1)},\dots,\mathbf{Z}_{0}^{(n)}\bigr{)}\,,$ we recursively simulate $R+1$ child configurations at each time step $t$, yielding $\{\mathbf{Z}_{t+1}[r]\}_{r=0,\dots,R}$. We designate the branch $r=0$ as the main branch $\{\mathbf{Z}_{t}[0]\}_{t=0,\dots,T-1}$, from which we extract the conditioning tuples $\bigl{(}\mathbf{Z}_{t}^{(i)},\,\{\mathbf{Z}_{t}^{(j)}\}_{j\in\mathcal{N}_{t}^{(i)}}\bigr{)}$ for all agents $i$. The remaining $R$ sibling branches at each $t$ supply the target next states $\mathbf{Z}_{t+1}^{(i)}$, ensuring that each context yields multiple outcomes.
@@
12:    $\nabla_{\phi,\omega}||\epsilon-\epsilon_{\phi}(\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau),\boldsymbol{c}_{t}^{(i)})||^{2}$
@@
Learning procedure. Our framework uses these data sets to train the Graph Diffusion Network. It minimizes the expected denoising loss over the outcomes observed in the ramification data (see Algorithm 1). At each training iteration, it uniformly samples a time index $t$ and extracts the conditioning pair $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\})$ from the main branch $\mathbf{Z}_{t}[0]$. We compute the interaction embedding $\mathbf{g}_{t}^{(i)}$ via Equation (3), then draw a diffusion step $\tau$ to form the condition vector $\mathbf{c}_{t}^{(i)}$, and uniformly select one of the $R$ next-state realizations to obtain the target $\mathbf{Z}_{t+1}^{(i)}$. Finally, we minimize the denoising loss in Equation (5) by backpropagating through both the diffusion model and the GNN. More details about the architecture can be found in Supplementary Section A.

### v2

@@ Abstract
Agent-Based Models (ABMs) are powerful tools for studying emergent properties in complex systems. In ABMs, agent behaviors are governed by local interactions and stochastic rules. However, these rules are, in general, non-differentiable, which limits the use of gradient-based methods for optimization, and thus integration with real-world data. We propose a novel framework to learn a differentiable surrogate of any ABM by observing its generated data. Our method combines diffusion models to capture behavioral stochasticity and graph neural networks to model agent interactions. Distinct from prior surrogate approaches, our method introduces a fundamental shift: rather than approximating system-level outputs, it models individual agent behavior directly, preserving the decentralized, bottom-up dynamics that define ABMs. We validate our approach on two ABMs (Schelling’s segregation model and a Predator-Prey ecosystem) showing that it replicates individual-level patterns and accurately forecasts emergent dynamics beyond training. Our results demonstrate the potential of combining diffusion models and graph learning for data-driven ABM simulation.
@@ 1 Introduction
Agent-Based Models (ABMs) are computational frameworks in which autonomous “agents” interact with each other and their environment, leading to emergent collective behavior (wilensky2015introduction). ABMs are typically characterized by: (i) a well-defined network of interactions, where the state of each agent is influenced by the states of a specific set of other agents, usually from the previous time step; (ii) stochasticity, meaning that agents’ decisions incorporate a degree of randomness, producing probability distributions over multiple runs that capture real-world uncertainty and variation. ABMs have proven to be a powerful tool for developing and refining theoretical understanding, particularly in identifying minimal sets of micro-level rules that generate realistic macro-level outcomes (railsback2019agent). In this sense, they have been instrumental in modeling a diverse range of phenomena (castellano2009statistical), including structure formation in biological systems, pedestrian traffic, urban aggregation, and opinion dynamics. More recently, ABMs have demonstrated their value as forecasting tools (poledna2020economic), such as in predicting the economic impacts of the COVID-19 pandemic (pichler2022forecasting).
@@
However, this progress is occurring despite the absence of principled methods to systematically align ABMs with real-world data. While various approaches have been proposed for calibrating macro-level parameters of ABMs (platt2020comparison), there are still no established methods for tuning the micro-level behaviors of individual agents to match observed data. One potential approach is to manually construct a probabilistic model that replicates the ABM and then use its likelihood function to estimate individual state variables (monti2023learning). However, this method requires the manual development of an ad-hoc probabilistic framework that reproduces the original ABM. Thus, what is currently missing is a fully automated method for deriving a learnable, differentiable model directly from an ABM.
@@
Figure 1: Overview of the training and generation pipeline for differentiable surrogates of Agent-Based Models. The top-left panel illustrates the training process: we run simulations using the original ABM, and use the resulting trajectories to train the differentiable surrogate. The top-right panel shows the structure of the ABM-generated data using the Predator-Prey model as an example: a state at time step t1t{-}1 gives rise to multiple possible states at time $t$, one of which is chosen to generate further possible states at t+1t{+}1. Colored cells highlight the behavior of a specific “prey” agent — green for “move,” red for “die,” and pink for “reproduce.” The bottom panel shows the generation phase: given a new observed state, the trained surrogate simulates plausible future states.
@@
In this work, we propose a novel approach to address this challenge: combining a graph neural network with a diffusion model to learn a differentiable surrogate of an ABM, from ABM-generated data. We refer to this method as a Graph Diffusion Network (GDN). In this framework, a graph neural network captures the interactions that govern the evolution of each agent’s state in response to other agents, while the diffusion model learns the distribution of possible state transitions, conditioned on these interactions. A central aspect of our approach is its explicit modeling of individual agent behavior. Rather than treating the system as a whole, we focus on how each agent acts as an independent entity, while also incorporating the influence of other agents on its decisions. This approach ensures that the emergent dynamics remain faithful to the decentralized nature of ABMs. By constraining the surrogate to model only micro-level rules, it cannot rely on shortcuts that predict only macro-level outcomes, preserving the distributed logic of the original ABM.
@@
Our approach draws inspiration from previous work on using neural network models to emulate deterministic cellular automata (grattarola2021learning). However, we extend this idea to the broader domain of ABMs by introducing a crucial component: stochasticity. By incorporating stochasticity, our architecture can learn directly from ABM-generated data traces, making it adaptable to a wide variety of agent-based models across diverse real-world applications. Furthermore, since our method is trained on data traces, it can seamlessly integrate empirical observations alongside simulated data, thus being potentially applicable to real-world scenarios. In this sense, our work represents a first step toward developing a comprehensive methodology for creating easy-to-use, learnable ABMs.
@@ 2 Background
To illustrate this structure, we consider two ABMs used throughout the paper. The first is the well-known model by Schelling schelling1971dynamic, where $\mathbf{Z}_{t}$ captures agents’ positions and colors, and $\Theta$ indicates preference for same-color neighbors. Even with some tolerance for neighbors of different colors, agents often form segregated clusters (ubarevivciene2024fifty). This clear mismatch between individual preferences and aggregate outcomes is a classic example of emergence. The second model is a predator-prey model (tang2022data; wilensky2015introduction), describing the ecological dynamics between two interacting species, with one acting as predator and the other as prey, similarly to the Lotka-Volterra equations. In this ABM, $\mathbf{Z}_{t}$ includes agent position and type (prey-predator), while $\Theta$ governs the probability to move, reproduce, or die. This model replicates the cyclical predator-prey population dynamics, typical of Lotka-Volterra systems, while also capturing complex spatial patterns reminiscent of spatial evolutionary games (nowak1992evolutionary). Both the Schelling and predator-prey models are widely recognized as canonical ABMs and are frequently used as testbeds for the development of novel calibration and surrogate techniques  Squazzoni2010; Jamali2024; McLane2011; tang2022data; TenBroeke2021; Murphy2020.
@@
ABMs have traditionally been powerful for theory generation, but in recent years, they have become increasingly data-driven (pangallo2024datadriven). To align ABM output with empirical data, most efforts focus on calibrating parameters $\Theta$ so that model-generated summary statistics match observed ones (platt2020comparison; quera2023blackbirds). Less attention, however, has been paid to estimating agent states $\mathbf{Z}_{t}$, which is key for matching time series further to summary statistics. Some researchers use data assimilation methods like particle filters (lux2018estimation) or ensemble Kalman filters (oswald2025agent) to infer $\mathbf{Z}_{t}$. A more principled alternative is to make ABMs differentiable, enabling the maximization of a likelihood function via gradient descent and automatic differentiation (monti2020learning; monti2023learning). While differentiability is straightforward for simple stochastic behaviors, such as those governed by Bernoulli trials (arya2022automatic), it becomes far more challenging for complex behaviors like those observed in the Schelling and predator-prey models.
@@
To address this and other challenges in ABMs, researchers have increasingly turned to more tractable surrogates, also known as meta-models or emulators (kennedy2001bayesian; farah2014bayesian; glielmo2023reinforcement). Surrogate models typically learn directly the mapping from parameters $\Theta$ to static summary statistics, disregarding individual behavior and model dynamics. For instance, a surrogate in Lamperti et al. lamperti2018agent maps $\Theta$ to the mean growth rate of the economy. More recent research has also explored the emulation of model dynamics. Grattarola et al. grattarola2021learning use Graph Neural Networks to approximate cellular automata, which can be seen as a special case of ABMs with deterministic interaction rules. Dyer et al. dyer2024interventionally propose Ordinary Differential Equation emulators to construct interventionally consistent surrogates, ensuring that micro-state interventions produce results aligned with macro-state interventions. Casert et al. casert2024learning employ Transformers to model the transition of physical systems from one configuration to another, in terms of their transition rates rather than reproducing individual agent behavior. Their method is tailored to physical systems in continuous time, making it inapplicable for interacting agents in the general case, since it requires explicitly enumerating allowed transitions between configurations.
@@
To achieve this goal, our framework relies on a novel combination of graph neural networks and diffusion models. Diffusion models (ddpm_original_paper) were first introduced in the context of image generation, where they demonstrated impressive generation capabilities (diffusion_vision_review), and were then applied to other domains (tabddpm). A number of works addressed graph data (graph_diffusion_review), for example in molecule modeling (molecule_modeling) and protein structure generation (protein_diffusion). However, these works focus on the generation of graphs, while our architecture learns to generate random samples that are conditioned on information found on a graph. To the best of our knowledge, our work is the first application of this generative framework to individual behavior modeling in simulation systems, such as ABMs.
@@ 3 Methods
To effectively model these components in the same individual-level view, we leverage respectively (i) message‐passing GNNs (message_passing_gnn), which model the relationship between the evolution of an agent’s state and the state of its neighbors on the graph; (ii) conditional diffusion models (conditionalimagesynthesisdiffusionreview), generative architectures well-suited to learning complex, multimodal distributions, allowing us to capture the intrinsic stochasticity of agent behavior.
@@
Overview. To learn the distribution $P_{\Theta}$, the training phase requires observations of different outcomes given the same starting conditions. To do so, in our framework, we use the original ABM to generate a data set as a ramification of possible states, namely $\bigl(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}\bigr)\;\longrightarrow\;\mathbf{Z}_{t+1}^{(i)}$ (see Figure 1). Our Graph Diffusion Network then approximates the stochastic kernel $P_{\Theta}$ by integrating a Message‐Passing GNN with a Conditional Diffusion Model, of learnable parameters $\omega$ and $\phi$ respectively. The GNN aggregates each agent’s state $\mathbf{Z}_{t}^{(i)}$ and its neighbors’ states $\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}$ via permutation‐invariant message and readout functions to produce an interaction embedding $\mathbf{g}_{t}^{(i)}$. This embedding acts as a compact representation of the information coming from $i$’s neighbors at $t$, affecting the distribution of possible states of agent $i$ at time $t+1$. As such, it is passed to the diffusion model: conditioned on $\mathbf{Z}_{t}^{(i)}$ and $\mathbf{g}_{t}^{(i)}$, the diffusion model learns to transform a sample of Gaussian noise into a possible instance of the next state $\mathbf{Z}_{t+1}^{(i)}$~. By minimizing the standard denoising loss over all observed transitions, this hybrid architecture captures both the graph‐structured interactions and the inherent stochasticity of agent dynamics. The trained model $GDN_{\phi,\omega}$ is therefore able to generate, given a graph $G_{t}$ of interacting agents and the state of each one $\mathbf{Z}^{(i)}_{t}$, a sequence of possible next states $\mathbf{Z}^{(i)}_{t+1}$. The consecutive application of $GDN_{\phi,\omega}$ allows for reproducing the behavior of the original model. We now describe in detail each of these components.
@@
Message‐passing GNN. The GNN operates on the provided interaction graph $G_{t}=(A,E_{t})$, that we assume to be known or to be computable from $\mathbf{Z}_{t}$ (e.g., in the Schelling model, the position of the agents determines who interacts with whom). For each agent $i$, the GNN aggregates its state $\mathbf{Z}_{t}^{(i)}$ together with each neighbor’s state $\mathbf{Z}_{t}^{(j)}$ via a permutation‐invariant operator $\bigoplus$, and then feeds the concatenated result through an MLP $f_{\omega}$ (message_passing_gnn), that is $\mathbf{g}_{t}^{(i)}=f_{\omega}\!\Bigl(\mathbf{Z}_{t}^{(i)},\,\bigoplus_{j\in N_{t}^{(i)}}\bigl(\mathbf{Z}_{t}^{(i)},\mathbf{Z}_{t}^{(j)}\bigr)\Bigr)$. The resulting vector $\mathbf{g}_{t}^{(i)}$ captures how $i$’s local neighborhood influences its next‐state distribution. In practice, for the operator $\bigoplus$, we used sum aggregation for Predator-Prey and mean aggregation for Schelling, since the latter’s dynamics depend on the degree. Minimal experimental evaluation can guide practitioners toward the most suitable operator.
@@
Conditional diffusion model. The diffusion model then learns the distribution over future states given this output from the graph and the state of a given agent. Diffusion models do so by reversing a fixed Gaussian noising process (ddpm_original_paper). The obtained denoising process, indexed by $\tau\in\{\tau_{\text{max}},\dots,0\}$, starts from a sample of Gaussian noise $\mathbf{x}_{\tau_{max}}\sim\mathcal{N}(\mathbf{0,I})$ and in a sequence of denoising diffusion steps transforms it into a possible next state $\mathbf{x}_{0}\approx\mathbf{Z}_{t+1}^{(i)}$. In this setting, we denote the general latent $\mathbf{x}_{\tau}$ as $\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau)$. Each step of this process receives as input (i) the agent’s current state $\mathbf{Z}_{t}^{(i)}$, (ii) its interaction embedding $\mathbf{g}_{t}^{(i)}$, and (iii) a sinusoidal positional embedding of $\tau$. These inputs are first transformed by MLPs to form the condition vector $\mathbf{c}_{t}^{(i)}$. Then, a feed‐forward network $\phi$ is trained to predict the noise residual $\epsilon_{\phi}$, i.e., the change to apply to the input $\tilde{\mathbf{Z}}_{t+1}^{(i)}(\tau)$ to continue the denoising process.
@@
Ramification data set. Given these two components, our framework uses the original ABM to produce a ramifications data set (see Figure 1). Such data set follows one main branch that specifies the evolution of the ABM, and multiple alternative stochastic evolutions of each time step from time $t$ to time $t+1$. This method makes it possible to expose the model to multiple stochastic successors from identical conditioning context, while avoiding exponential growth in the number of histories. Starting from an initial configuration $\mathbf{Z}_{0}=\bigl(\mathbf{Z}_{0}^{(1)},\dots,\mathbf{Z}_{0}^{(n)}\bigr)\,$, we recursively simulate $R+1$ child configurations at each time step $t$, yielding $\{\mathbf{Z}_{t+1}[r]\}_{r=0,\dots,R}$. We designate the branch $r=0$ as the main branch $\{\mathbf{Z}_{t}[0]\}_{t=0,\dots,T-1}$, from which we extract the conditioning tuples $\bigl(\mathbf{Z}_{t}^{(i)},\,\{\mathbf{Z}_{t}^{(j)}\}_{j\in N_{t}^{(i)}}\bigr)$ for all agents $i$. The remaining $R$ sibling branches at each $t$ supply the target next states $\mathbf{Z}_{t+1}^{(i)}$, ensuring that each context yields multiple outcomes.
@@
[…]
@@
Learning procedure. Our framework uses these data sets to train the Graph Diffusion Network. It minimizes the expected denoising loss over the outcomes observed in the ramification data (see Algorithm 1). At each training iteration, it uniformly samples a time index $t$ and extracts the conditioning pair $(\mathbf{Z}_{t}^{(i)},\{\mathbf{Z}_{t}^{(j)}\})$ from the main branch $\mathbf{Z}_{t}[0]$. We compute the interaction embedding $\mathbf{g}_{t}^{(i)}$ via Equation (3), then draw a diffusion step $\tau$ to form the condition vector $\mathbf{c}_{t}^{(i)}$, and uniformly select one of the $R$ next‐state realizations to obtain the target $\mathbf{Z}_{t+1}^{(i)}$. Finally, we minimize the denoising loss in Equation (5) by backpropagating through both the diffusion model and the GNN. More details about the architecture, and a discussion on computational costs and scalability can be found in Supplementary Section A.